Drive the cost function for the following production function

Q=min(2*X1+X2, X1+2*X2)
Answer #1

The general cost function is

TC= w.X1+r.X2

Where w is cost of X1 and r is cost of X2

At tangency condition for the given production function

2X1+X2= X1+2X2

X1=X2

C= (w+r)X1 or (w+r)X2